# NumGrad-Pull: Numerical Gradient Guided Tri-plane Representation for Surface Reconstruction from Point Clouds

### v1

@@ 4.4 Ablation Studies
4.4.1 Effect of Tri-plane Resolution
@@
Considering that tri-plane resolution impacts surface reconstruction quality, we empirically evaluate the performance of NumGrad-Pull using various resolutions, 8, 16, 32, and 64, on the FAMOUS dataset. Visual results are shown in Fig. 4, and quantitative results are provided in Tab. 4. Our findings indicate that as tri-plane resolution increases, surface reconstruction quality (measured by $CD_{\ell_{2}}$) improves. The most substantial gain is observed when increasing the resolution from 8 to 16, while increasing to 32 and 64 resolution yields marginal improvements. The significant quality boost from 8 to 16 suggests that an $8\times 8$ tri-plane lacks sufficient capacity to store the necessary information for accurate surface reconstruction, as shown in Fig. 4(b). Higher resolutions enhance the network’s ability to capture fine structures; however, with a $64\times 64$ tri-plane, while the metric show slight improvements, the visual results reveal unwanted surface unevenness, indicating that excessive high-frequency noise is being captured. Overall, to balance reconstruction quality with robustness against noise, we apply a $32\times 32$ tri-plane for our main experiments.
@@
We also present efficiency comparisons between recent methods with our NumGrad-Pull using different tri-plane resolutions in Tab. 4. Since our proposed tri-plane-based neural SDF can significantly improve query speed, we achieve a 1.8× speedup over Neural-Pull [27] and a 5.4× speedup over IF [22], while retaining superior surface reconstruction quality. Furthermore, the tri-plane representation exhibits excellent scalability with respect to resolution: as the tri-plane resolution increases, the query speed remains stable. This stability is attributed to the feature interpolation process being an $O(1)$ operation, ensuring that query speed is unaffected by resolution changes.
4.4.2 Effect of Each Module
We conduct a detailed ablation study to validate the effectiveness of each proposed module, including the following model variants: A) full model without our data sampling strategy; B) full model without the tri-plane structure; C) full model without numerical gradients; and D) full model without progressive tri-plane expansion. These variants are compared with a baseline model Neural-Pull [27] and our complete model incorporating all proposed modules. We evaluate the performances on the FAMOUS dataset, provide quantitative results in Tab. 5, and present a visualization of the hand sample in Fig. 5.
@@ 4.4.2 Effect of Each Module
Our full model achieves the best visual and quantitative results, capturing all essential details without artifacts. In contrast, each variant shows varying degrees of performance degradation. Model A lacks a comprehensive data sampling strategy, sampling query points only near the object surface, similar to Neural-Pull. This model suffers from insufficient supervision for regions far from the surface, leading to poorly trained tri-plane features at distant positions and causing random predictions and artifacts. Model B replaces the tri-plane with an 8-layer fully connected network, achieving results close to our full model but with reduced reconstruction detail and training stability. While it retains more structural detail than the baseline by employing numerical gradients, it is approximately 3.2 times slower due to the slower query speed of the fully connected network compared to the tri-plane structure. Model C fails to train effectively due to the adoption of analytical gradients, which only propagate supervision to local grids and causes instability and corrupted outputs. Model D struggles without using our proposed progressive expansion strategy, which is crucial for preventing the network from getting stuck in local optima. This experimental comparison highlights the necessity of our strategy for the effective training of the tri-plane structure. Overall, these results demonstrate that each of the proposed modules is essential, contributing to the superior performance of the full model.
@@
Table 5: Effect of each proposed module of our method on surface reconstruction for the FAMOUS dataset, measured by $CD_{\ell_{2}}$, scaled by $10^{4}$.
@@
5 Future Work
While our proposed NumGrad-Pull shows promising results, there are limitations to address in future work. In our ablation study on tri-plane resolution, our method captures high-frequency noise at higher resolutions, limiting its effectiveness in finer-grained modeling and highlighting a potential improvement area. Additionally, we plan to explore various regularization techniques to improve the robustness of surface reconstruction against noisy and sparse point clouds. Moreover, although our method is primarily introduced for object-level surface reconstruction, the tri-plane structure holds strong potential for scene-level modeling. We intend to conduct a more comprehensive evaluation using scene-level reconstruction data to further validate and extend the applicability of our approach.
6 Conclusion
In this paper, we introduce NumGrad-Pull, a novel approach for high-fidelity surface reconstruction from unoriented point clouds. Our method employs a hybrid explicit–implicit tri-plane representation to improve query speed and enhance reconstruction fidelity. Furthermore, we propose a numerical gradients-based technique to stabilize the training process and a progressive expansion strategy to fully leverage the capabilities of the tri-plane structure. Extensive experiments demonstrate that our NumGrad-Pull consistently outperforms state-of-the-art methods across various challenging scenarios, showcasing significant effectiveness in surface reconstruction problems.
Broader Impact: We believe NumGrad-Pull has strong potential to improve 3D content creation and streamline the conversion of raw 3D data into production-ready formats. While we do not foresee any immediate harmful uses, we emphasize the importance of users exercising caution to minimize potential risks, given that surface reconstruction tools could also be leveraged for malicious purposes.

### v2

@@ IV-E Ablation Studies
TABLE IV: Effect of different tri-plane configurations on surface reconstruction for the FAMOUS dataset [10] compared with recent methods, measured by $CD_{\ell_{2}}$ and scaled by $10^{4}$.
Metric 8 16 32 64 NP [24] IF [19] $CD_{\ell_{2}}$ $\downarrow$ 9.31 0.45 0.39 0.35 1.00 0.80 Speed (iter/s) $\uparrow$ 219 231 227 225 122 42
IV-E1 Effect of Tri-plane Resolution
@@
Considering that tri-plane resolution impacts surface reconstruction quality, we empirically evaluate the performance of NumGrad-Pull using various resolutions, 8, 16, 32, and 64, on the FAMOUS dataset. Visual results are shown in Fig. 5, and quantitative results are provided in Tab. IV. Our findings indicate that as tri-plane resolution increases, surface reconstruction quality (measured by $CD_{\ell_{2}}$) improves. The most substantial gain is observed when increasing the resolution from 8 to 16, while increasing to 32 and 64 resolution yields marginal improvements. The significant quality boost from 8 to 16 suggests that an $8\times 8$ tri-plane lacks sufficient capacity to store the necessary information for accurate surface reconstruction, as shown in Fig. 5(b). Higher resolutions enhance the network’s ability to capture fine structures; however, with a $64\times 64$ tri-plane, while the metric show slight improvements, the visual results reveal unwanted surface unevenness, indicating that excessive high-frequency noise is being captured. Overall, to balance reconstruction quality with robustness against noise, we apply a $32\times 32$ tri-plane for our main experiments.
@@
We also present efficiency comparisons between recent methods with our NumGrad-Pull using different tri-plane resolutions in Tab. IV. Since our proposed tri-plane-based neural SDF can significantly improve query speed, we achieve a 1.8× speedup over Neural-Pull [24] and a 5.4× speedup over IF [19], while retaining superior surface reconstruction quality. Furthermore, the tri-plane representation exhibits excellent scalability with respect to resolution: as the tri-plane resolution increases, the query speed remains stable. This stability is attributed to the feature interpolation process being an $O(1)$ operation, ensuring that query speed is unaffected by resolution changes.
IV-E2 Effect of Query Ratio
To further analyze the efficiency trade-off of our tri-plane formulation, we evaluate different values of the query ratio, ranging from $1/2$ to $1/32$, which determine the fraction of query points used to train the tri-plane-based SDF representation in Eq. 2.
TABLE V: Effect of different query point ratios used to train the Tri-plane-based SDF representation.
Ratio ABC [15] FAMOUS–3DBenchy [10] ShapeNet–Sofa [6] SRB–Gargoyle [34] 1/2 0.11 0.08 0.024 0.044 1/4 0.11 0.05 0.021 0.041 1/8 0.09 0.04 0.020 0.033 1/16 0.12 0.04 0.022 0.037 1/32 0.14 0.09 0.029 0.041
We report the quantitative results on five representative datasets in Tab. V. Specifically, the ratio of $1/8$ achieves the best overall performance across all evaluated cases. When the ratio is too large (e.g., $1/2$ or $1/4$), the model tends to rely excessively on dense tri-plane queries, which increases sensitivity to high-frequency noise and does not consistently improve reconstruction quality. Conversely, when the ratio is too small (e.g., $1/16$ or $1/32$), the supervision and feature representation of the tri-plane branch become insufficient, leading to degraded reconstruction quality, as evidenced by the noticeable performance drop on ABC and SRB samples. The results suggest that a moderate query ratio can effectively balance the benefits of fast, robust, and generalizable tri-plane-based SDF modeling. We thus adopt the $1/8$ ratio as the default setting in our main experiments.
IV-E3 Effect of Each Module
@@
We conduct a detailed ablation study to validate the effectiveness of each proposed module, including the following model variants: A) full model without our data sampling strategy; B) full model without the tri-plane structure; C) full model without numerical gradients; and D) full model without progressive tri-plane expansion. These variants are compared with a baseline model Neural-Pull [24] and our complete model incorporating all proposed modules. We evaluate the performances on the FAMOUS dataset, provide quantitative results in Tab. VI, and present a visualization of the hand sample in Fig. 6. Our full model achieves the best visual and quantitative results, capturing all essential details without artifacts. In contrast, each variant shows varying degrees of performance degradation. Model A lacks a comprehensive data sampling strategy, sampling query points only near the object surface, similar to Neural-Pull. This model suffers from insufficient supervision for regions far from the surface, leading to poorly trained tri-plane features at distant positions and causing random predictions and artifacts. Model B replaces the tri-plane with an 8-layer fully connected network, achieving results close to our full model but with reduced reconstruction detail and training stability. While it retains more structural detail than the baseline by employing numerical gradients, it is approximately 3.2 times slower due to the slower query speed of the fully connected network compared to the tri-plane structure. Model C fails to train effectively due to the adoption of analytical gradients, which only propagate supervision to local grids and causes instability and corrupted outputs. Model D struggles without using our proposed progressive expansion strategy, which is crucial for preventing the network from getting stuck in local optima. This experimental comparison highlights the necessity of our strategy for the effective training of the tri-plane structure. Overall, these results demonstrate that each of the proposed modules is essential, contributing to the superior performance of the full model.
@@
TABLE VI: Effect of each proposed module of our method on surface reconstruction for the FAMOUS dataset, measured by $CD_{\ell_{2}}$, scaled by $10^{4}$.
@@
V Conclusion
In this paper, we introduce NumGrad-Pull, a novel approach for high-fidelity surface reconstruction from unoriented point clouds. Our method employs a hybrid explicit–implicit tri-plane representation to improve query speed and enhance reconstruction fidelity. Furthermore, we propose a numerical gradients-based technique to stabilize the training process and a progressive expansion strategy to fully leverage the capabilities of the tri-plane structure. Extensive experiments demonstrate that our NumGrad-Pull consistently outperforms state-of-the-art methods across various challenging scenarios, showcasing significant effectiveness in surface reconstruction problems. We believe NumGrad-Pull has strong potential to improve 3D content creation and streamline the conversion of raw 3D data into production-ready formats. While we do not foresee any immediate harmful uses, we emphasize the importance of users exercising caution to minimize potential risks, given that surface reconstruction tools could also be leveraged for malicious purposes.